A system comprising blocks, a light frictionless pulley, a frictionless incline at 30°, and a rope connecting 6.0-kg and 4.0-kg blocks on the incline with the 6.0-kg block being lower than the 4.0-kg block. The rope extends to a 9.0-kg block. The 9.0-kg block hangs over the pulley on the side of the incline and accelerates downward when the system is released from rest. The tension in the rope connecting the 6.0-kg block and the 4.0-kg block is closest to


  • A. 30 N.

 

  •  B. 39 N.

 

  •  C.  36 N.


  •  D. 42 N.
1
Expert's answer
2021-07-01T04:24:01-0400

Gives


"m_1=6,m_2=4kg,M=9kg,\theta=30°"


Find acceleration (a)


"(m_2+m_1)a-Ma =Mg-m_2gsin\theta -m_1gsin\theta"

"a=\frac{Mg-(m_2+m_1)gsin\theta}{m_1+m_2-M}"

Put value

"a=\frac{9g-(6+4)gsin\theta}{6+4-9}"

"a=\frac{9g-(10)gsin30°}{1}=4g"

"g=9.8m\/sec^2"

Tension


"2T=(m_1+m_2)gcos30\u00b0=10gcos30\u00b0=84N"

T=42N


Option (d) is correct option